Given an isosceles right triangle with sides length 5 cm, solve for the hypotenuse
Answer by John10(297)

We know that the isosceles right triangle has two equal sides and a 90 degree angle between them. And we know that each side is equal to 5. so we will find the hypotenuse side by the Pythagorean Theorem.
Hint: Draw a picture will be better to see.
Pythagorean theorem: h^2 = a^2 + b^2 = 5^2 + 5^2 = 25 + 25 = 50 so h = sqrt(50) or h = 5sqrt(2)
Hope it helps:)
John10